The ratio of number of pupils who visited the museum to the number of pupils who visited the zoo is 9:4. If 45 more pupils visited the museum than the zoo. How many pupils visited the zoo?
Answer by ikleyn(52787)

.

The number of pupils who visited the museum is 9x;  the number of pupils who visited the zoo is 4x,


where x is the common measure of these two quantities.


We are given that  9x - 4x = 45.   It implies  5x = 45,  i.e.  x = 45/5 = 9.


Hence, the number of pupils visiting zoo is  4x = 4*9 = 36.   ANSWER